The equation of common tangent of the parabolas y^2=32x and x^2=108y

Grade:12th pass

1 Answers

ashu
8 years ago
first write the equation of tan gent in slope form for both the parabolas
 i.e
y=mx+a/m       -1                 and                 x=my+a/m                     -2
now compare the cofficients of x,y,m,a
get the value of m and put in
and get the answer which is 2x - 3y=36